# MIRROR: Multi-Modal Pathological Self-Supervised Representation Learning via Modality Alignment and Retention

### v1

@@ Abstract
Histopathology and transcriptomics are fundamental modalities in cancer diagnostics, encapsulating the morphological and molecular characteristics of the disease. Multi-modal self-supervised learning has demonstrated remarkable potential in learning pathological representations by integrating diverse data sources. Conventional multi-modal integration methods primarily emphasize modality alignment, while paying insufficient attention to retaining the modality-specific intrinsic structures. However, unlike conventional scenarios where multi-modal inputs often share highly overlapping features, histopathology and transcriptomics exhibit pronounced heterogeneity, offering orthogonal yet complementary insights. Histopathology data provides morphological and spatial context, elucidating tissue architecture and cellular topology, whereas transcriptomics data delineates molecular signatures through quantifying gene expression patterns. This inherent disparity introduces a major challenge in aligning these modalities while maintaining modality-specific fidelity. To address these challenges, we present MIRROR, a novel multi-modal representation learning framework designed to foster both modality alignment and retention. MIRROR employs dedicated encoders to extract comprehensive feature representations for each modality, which is further complemented by a modality alignment module to achieve seamless integration between phenotype patterns and molecular profiles. Furthermore, a modality retention module safeguards unique attributes from each modality, while a style clustering module mitigates redundancy and enhances disease-relevant information by modeling and aligning consistent pathological signatures within a clustering space. Extensive evaluations on The Cancer Genome Atlas (TCGA) cohorts for cancer subtyping and survival analysis highlight MIRROR’s superior performance, demonstrating its effectiveness in constructing comprehensive oncological feature representations and benefiting the cancer diagnosis.
@@ I Introduction
Histopathology images are widely regarded as the gold standard in cancer diagnosis, offering critical insights into the presence, type, grade, and prognosis of cancer [1]. These images encapsulate a wealth of morphological features that serve as the foundation of cancer diagnostics [2, 3, 4, 5]. Meanwhile, advancements in high-throughput sequencing technologies, such as polymerase chain reaction (PCR) [6], have further expanded oncological diagnostic capabilities by enabling the analysis of molecular data, including transcriptomics data, which delineates gene expression profiles, offering molecular signatures of the disease. The integration of the morphological and molecular modalities within multi-modal diagnostics significantly enhances the accuracy of cancer diagnosis and prognosis prediction, providing a more comprehensive and precise understanding of the disease.
@@ II Related Work
II-A Vision Self-Supervised Learning
Recent advancements in SSL in computer vision (CV) have significantly reduced the reliance on labeled data while achieving performance comparable to, or even surpassing, that of supervised methods. SSL can be categorized into three main styles: contrastive learning, generative learning, and hybrid methods that integrate both approaches. Contrastive learning models [17, 18] achieve representation learning by drawing closer the embeddings of the same image under different augmentations while separating those of different images. Generative approaches [19, 20] aim to reconstruct or predict missing or unobserved portions of the input using the information from the observed parts. Hybrid methods, such as [21, 22], integrate the strengths of contrastive and generative learning to simultaneously capture both discriminative and semantic information effectively.
@@ II-C Histopathology and Transcriptomics Multi-Modal Learning
Multi-modal integration methods [12, 43, 44, 11, 45, 14, 46] that combine histopathology and transcriptomics have shown impressive capabilities in various pathological tasks, including cancer diagnosis and survival analysis. While these modalities share some information, they also contain a vast amount of comprehensive and distinct insights into the disease, reflecting their inherent heterogeneity. Despite improvements in performance, existing multi-modal methods predominantly focus on integration and alignment, often overlooking the rich and diverse information embedded in the unique characteristics of each modality.
@@ III-A2 RNA Encoder
Figure 3: Transcriptomics data distributions in the TCGA-NSCLC dataset. The top panel presents a heatmap visualization of transcriptomics data from two subtypes in TCGA-NSCLC, namely TCGA-LUAD and TCGA-LUSC. The data distribution exhibits substantial variability and clear subtype distinction after preprocessing, providing a robust foundation for representation learning. The bottom bar plot highlights the top 10 most variant genes in the TCGA-NSCLC dataset, identified with reference to the COSMIC database, demonstrating extraordinary biological explainability.
@@ III-B Preliminaries
As discussed in Section I, the output of each encoder can be conceptually decomposed into four parts: (1) disease-relevant, modality-shared; (2) disease-relevant, modality-specific; (3) disease-irrelevant, modality-shared; (4) disease-irrelevant, modality-specific. Hence, for the $i$-th paired sample, the outputs of the slide and RNA encoders can be written as
@@
where “r” (relevant) vs. “i” (irrelevant) indicates disease relevance, and “s” (shared) vs. “u” (unique) indicates modality traits. While we speak of disease relevance for clarity, in practice these relevant subspaces can capture any shared pathological characteristics that drive meaningful variation in the data. Additionally, for notational conciseness, $\mathbf{S}$ and $\mathbf{S}^{\textit{[CLS]}}$ are treated equivalently in this context.
@@ III-C Modality Alignment Module
$\mathbf{S}_{\text{align}}^{i}\approx\mathbf{T}_{\text{align}}^{i}\quad\text{and}\quad\mathbf{S}_{\text{align}}^{i}\cdot\mathbf{T}_{\text{align}}^{j}\rightarrow 0\quad\forall i\neq j.$ (9)
where the dot product measures similarity, ensuring that only corresponding pairs exhibit high alignment while unrelated pairs are orthogonal in the shared latent space.
@@
III-D Modality Retention Module
$\mathbf{S}_{\text{retention}}=S_{\mathrm{r},u}+S_{\mathrm{i},u}.$ (11)
$\mathbf{T}_{\text{retention}}=T_{\mathrm{r},u}+T_{\mathrm{i},u}.$ (13)
TABLE I: Cancer subtype classification on TCGA-BRCA and TCGA-NSCLC.
@@ III-D Modality Retention Module
TABLE II: Cancer survival analysis on TCGA-BRCA and TCGA-NSCLC datasets.
@@
III-E Style Clustering Module
To facilitate the alignment of disease-relevant features across modalities, MIRROR introduces a set of learnable cluster centers, $\mathbf{P}\in\mathbb{R}^{D}$, which are normalized to prevent degeneracy. Each latent representation $z_{S}$ and $z_{T}$ is softly assigned to these centers, producing assignment distributions $\mathbf{S}_{\text{cluster}}$ and $\mathbf{T}_{\text{cluster}}$, respectively. The alignment between these assignments is encouraged through a bidirectional KL divergence penalty:
ensuring mutual consistency between the two modalities.
Formally, let $z_{S}$ and $z_{T}$ be $d$-dimensional latent vectors such that:
$L_{\text{style clustering}}=L_{\text{style}}+L_{\text{cluster}},$ (18)
@@ III-F Global Optimization Objective
$L=L_{\text{align}}+L_{\text{retention}}+L_{\text{style clustering}}.$ (19)
Minimizing this composite loss function promotes the enrichment of disease-relevant information, capturing both modality-shared and modality-specific components, denoted by $S_{r,s}$, $S_{r,u}$, $T_{r,s}$, and $T_{r,u}$. Simultaneously, it suppresses disease-irrelevant information, represented by $S_{i,s}$, $S_{i,u}$, $T_{i,s}$, and $T_{i,u}$. This selective optimization ensures that the learned feature representations concentrate on patterns essential for accurate disease characterization across modalities, thereby improving the model’s diagnostic and prognostic effectiveness.
Figure 4: Visualization of slide encoder attention weights on TCGA-BRCA and TCGA-NSCLC. Regions exhibiting higher attention scores predominantly correspond to malignant, tumor-bearing tissue, whereas areas with lower scores typically indicate normal regions.
@@ IV-A Datasets
Two cohorts from the publicly available TCGA dataset [16] were used for evaluation: Breast Invasive Carcinoma (BRCA) and Non-Small Cell Lung Cancer (NSCLC). The TCGA-BRCA cohort comprises two primary subtypes: Invasive Ductal Carcinoma (IDC) with 787 samples and Invasive Lobular Carcinoma (ILC) with 168 samples, totaling 955 samples. Similarly, the TCGA-NSCLC cohort includes two subtypes: Lung Adenocarcinoma (LUAD) with 541 samples and Lung Squamous Cell Carcinoma (LUSC) with 512 samples, totaling 1053 samples. Whole Slide Images (WSIs) from both cohorts were processed into patches at 20×20$\times$20 × magnification, and feature representations were extracted using ResNet-50 [52] and Phikon [53]. Corresponding transcriptomics data were collected from Xena [54] and preprocessed using the proposed novel pipeline, reducing the original 198,620 genes to 40,146 for the BRCA cohort and 10,234 for the NSCLC cohort, forming the novel refined transcriptomics datasets.
@@ IV-B1 Data Preprocessing
For WSIs, adhering to the conventional preprocessing protocol established in [55], the foreground tissue regions are first segmented using Otsu’s method [56]. These segmented regions are then divided into patches, forming an instance bag that is processed by a pre-trained patch encoder to extract feature representations. To enable batched training, we perform random sampling to select a fixed number of feature representations, employing replacement if the available number of representations is less than the required count.
@@ IV-C Downstream Tasks
Two tasks are used for downstream evaluations, namely subtype classification and survival analysis. We compared the proposed MIRROR against ABMIL [59], PORPOISE [60], Linear Classifier, and TANGLE [11], utilizing two distinct backbones under both 10-shot and linear probing settings. Among these methods, ABMIL exclusively takes histopathology as input, whereas the other models integrate both histopathology and transcriptomics data. To ensure fair comparisons and optimal performance, all methods are trained on the proposed transcriptomic datasets. Notably, the original TANGLE framework evaluates only histopathology features. To comprehensively assess its performance and maintain methodological consistency, we extend TANGLE by incorporating both histopathology and transcriptomic features, aligning its input with that of MIRROR.
@@ IV-C1 Subtype Classification
As illustrated in Table I, the proposed MIRROR consistently achieves superior performance, demonstrating its ability to effectively integrate histopathology and transcriptomic data for enhanced classification accuracy. On the TCGA-NSCLC dataset, MIRROR achieves the highest accuracy and F1 score across all settings, outperforming all baseline methods. Notably, its strong performance remains consistent regardless of the chosen backbone architecture, highlighting its robustness on the TCGA-NSCLC dataset. Similarly, on the TCGA-BRCA dataet, MIRROR achieves state-of-the-art results, surpassing the second best model’s (TANGLE) result by 1.5% in F1 score under linear probing setting with the Phikon backend. These findings validate the robustness and generalizability of MIRROR across different cancer subtypes and backbone architectures.
@@ IV-D Ablation Study
TABLE III: Ablation Study of Model Components on TCGA-BRCA.
@@
We conducted an ablation study to evaluate the effectiveness of the key components of MIRROR, including the modality alignment module, modality retention module, and style clustering module. The study was performed on the TCGA-BRCA dataset using linear probing with Phikon as the backbone, comparing MIRROR’s performance both with and without these components.
As shown in Table III, the results highlight the contributions of each module. The modality retention module, while leading to a slight decline in survival analysis performance, enhances subtype classification, demonstrating its ability to preserve modality-specific information. Additionally, the introduction of the style clustering module encourages the model to focus on disease-relevant patterns while mitigating redundancy, enabling the extraction of more informative features. This refinement is reflected in the improved feature retention performance observed during training. With all modules incorporated, MIRROR achieves the best overall performance, confirming the synergistic benefits of the proposed components.
IV-E Visualization
Figure 5: Visualization of transcriptomics features encoded by MIRROR on the TCGA-NSCLC dataset, compared to those produced by TANGLE. Orange dots represent samples from TCGA-LUSC, while blue dots correspond to TCGA-LUAD. MIRROR clearly yields more distinct and representative feature distributions.
@@ IV-E Visualization
As illustrated in Figure 4, we selected two representative samples from TCGA-BRCA and TCGA-NSCLC, respectively. The attention weights from the slide encoder are visualized using heatmaps to highlight regions of interest within the images. Areas shown in warmer colors indicate regions which the model pays more attention to. The results suggest MIRROR consistently attends to disease-relevant regions, particularly on regions identified as cancerous. This strong alignment showcases MIRROR’s extraordinary interpretability, confirming that it not only excels in downstream tasks but also provides transparent, biologically grounded decision-making insights.
Additionally, as depicted in Figure 5, we visualize features extracted by RNA encoders from TANGLE and MIRROR on the TCGA-NSCLC dataset. Orange and blue dots represent samples from TCGA-LUSC and TCGA-LUAD, respectively. Features generated by MIRROR exhibit significantly improved class separability, with minimal overlap between TCGA-LUSC and TCGA-LUAD clusters. In contrast, TANGLE’s feature space demonstrates considerable inter-class mixing, suggesting weaker discriminative capability. These findings highlight the advantage of MIRROR in generating more representative and structured feature distributions, ultimately contributing to superior model performance on downstream tasks.

### v2

@@ Abstract
Histopathology and transcriptomics are fundamental modalities in cancer diagnostics, encapsulating the morphological and molecular characteristics of the disease. Multi-modal self-supervised learning has demonstrated remarkable potential in learning pathological representations by integrating diverse data sources. Conventional multi-modal integration methods primarily emphasize modality alignment, while paying insufficient attention to retaining the modality-specific intrinsic structures. However, unlike conventional scenarios where multi-modal inputs often share highly overlapping features, histopathology and transcriptomics exhibit pronounced heterogeneity, offering orthogonal yet complementary insights. Histopathology data provides morphological and spatial context, elucidating tissue architecture and cellular topology, whereas transcriptomics data delineates molecular signatures through quantifying gene expression patterns. This inherent disparity introduces a major challenge in aligning these modalities while maintaining modality-specific fidelity. To address these challenges, we present MIRROR, a novel multi-modal representation learning framework designed to foster both modality alignment and retention. MIRROR employs dedicated encoders to extract comprehensive feature representations for each modality, which is further complemented by a modality alignment module to achieve seamless integration between phenotype patterns and molecular profiles. Furthermore, a modality retention module safeguards unique attributes from each modality, while a style clustering module mitigates redundancy and enhances disease-relevant information by modeling and aligning consistent pathological signatures within a clustering space. Extensive evaluations on The Cancer Genome Atlas (TCGA) cohorts for cancer subtyping and survival analysis highlight MIRROR’s superior performance, demonstrating its effectiveness in constructing comprehensive oncological feature representations and benefiting the cancer diagnosis. Code is available at https://github.com/TianyiFranklinWang/MIRROR.
@@ I Introduction
Histopathology images are widely regarded as the gold standard in cancer diagnosis, offering critical insights into the presence, type, grade, and prognosis of cancer [1]. These images encapsulate a wealth of morphological features that serve as the foundation of cancer diagnostics [2, 3, 4, 5, 6]. Meanwhile, advancements in high-throughput sequencing technologies, such as polymerase chain reaction (PCR) [7], have further expanded oncological diagnostic capabilities by enabling the analysis of molecular data, including transcriptomics data, which delineates gene expression profiles, offering molecular signatures of the disease. The integration of the morphological and molecular modalities within multi-modal diagnostics significantly enhances the accuracy of cancer diagnosis and prognosis prediction, providing a more comprehensive and precise understanding of the disease.
@@ II Related Work
II-A Self-Supervised Learning in Computer Vision
Recent advancements in SSL in computer vision (CV) have significantly reduced the reliance on labeled data while achieving performance comparable to, or even surpassing, that of supervised methods. SSL can be categorized into three main styles: contrastive learning [18, 19], generative learning [20, 21], and hybrid methods [22, 23] that integrate both approaches.
@@ II-C Histopathology and Transcriptomics Multi-Modal Learning
Multi-modal integration methods [13, 44, 45, 12, 46, 15, 47] that combine histopathology and transcriptomics have shown impressive capabilities in various pathological tasks, including cancer diagnosis and survival analysis.
In practice, these integration methods predominantly emphasize bridging these modalities through contrastive learning or alternative alignment techniques, such as clustering and cross-attention, to merge data from multiple sources into a unified feature space. Although these modalities share some information, they also contain a vast amount of comprehensive and distinct insights into the disease, reflecting their inherent heterogeneity. While recent advancements have yielded improved outcomes, many existing integration approaches overlook the rich and diverse insights embedded in the unique attributes of each modality.
@@ III-A2 RNA Encoder
Figure 3: Transcriptomics data distributions in the TCGA-NSCLC dataset. The top panel displays a heatmap visualization of transcriptomic data for two subtypes in TCGA-NSCLC: TCGA-LUAD on the left and TCGA-LUSC on the right. The data distribution exhibits substantial variability and clear subtype distinction after preprocessing, providing a robust foundation for representation learning. The bottom bar plot highlights the top 10 most variant genes in the TCGA-NSCLC dataset, identified with reference to the COSMIC database, demonstrating extraordinary biological explainability.
@@ III-B Preliminaries
As discussed in Section I, the output of each encoder can be conceptually decomposed into four parts: (1) disease-relevant, modality-shared; (2) disease-relevant, modality-specific; (3) disease-irrelevant, modality-shared; and (4) disease-irrelevant, modality-specific. Hence, for the $i$-th paired sample, the outputs of the slide and RNA encoders can be written as:
@@
where “$\mathrm{r}$” (relevant) vs. “$\mathrm{i}$” (irrelevant) indicates disease relevance, and “$s$” (shared) vs. “$u$” (unique) indicates modality traits. In this notation, the “$+$” denotes the composition of subspace features rather than simple vector addition. While we speak of disease relevance for clarity, in practice these relevant subspaces can capture any shared pathological characteristics that drive meaningful variation in the data. Additionally, for notational conciseness, $\mathbf{S}$ and $\mathbf{S}^{\textit{[CLS]}}$ are treated equivalently in this context.
@@ III-C Modality Alignment Module
$\mathbf{S}_{\text{align}}^{i}\approx\mathbf{T}_{\text{align}}^{i}\quad\text{and}\quad\mathbf{S}_{\text{align}}^{i}\cdot\mathbf{T}_{\text{align}}^{j}\rightarrow 0\quad\forall i\neq j,$ (9)
where the dot product measures similarity, ensuring that only corresponding pairs exhibit high alignment while unrelated pairs are nearly orthogonal in the shared latent space.
@@
TABLE I: Cancer subtype classification on TCGA-BRCA, TCGA-NSCLC, TCGA-RCC, and TCGA-COADREAD.
@@
TCGA- RCC ResNet-50 10-shot 0.752 0.629 0.937 0.846 0.752 0.673 0.938 0.832 0.942 0.853
±0.059 ±0.121 ±0.043 ±0.112 ±0.016 ±0.099 ±0.034 ±0.111 ±0.028 ±0.126
All Data 0.908 0.677 0.976 0.857 0.933 0.827 0.988 0.903 0.998 0.932
±0.017 ±0.144 ±0.031 ±0.194 ±0.007 ±0.116 ±0.007 ±0.138 ±0.003 ±0.150
Phikon 10-shot 0.839 0.733 0.951 0.854 0.828 0.727 0.950 0.867 0.952 0.931
±0.029 ±0.083 ±0.030 ±0.114 ±0.027 ±0.107 ±0.014 ±0.123 ±0.030 ±0.042
All Data 0.968 0.886 0.988 0.920 0.961 0.885 0.989 0.923 0.997 0.931
±0.015 ±0.133 ±0.011 ±0.015 ±0.011 ±0.130 ±0.001 ±0.147 ±0.003 ±0.149
TCGA- COADREAD ResNet-50 10-shot 0.815 0.542 0.813 0.448 0.819 0.502 0.831 0.539 0.834 0.556
±0.057 ±0.066 ±0.065 ±0.020 ±0.055 ±0.080 ±0.043 ±0.066 ±0.039 ±0.144
All Data 0.828 0.515 0.814 0.447 0.816 0.470 0.928 0.881 0.938 0.893
±0.051 ±0.091 ±0.065 ±0.020 ±0.069 ±0.064 ±0.042 ±0.051 ±0.022 ±0.017
Phikon 10-shot 0.659 0.495 0.813 0.448 0.644 0.510 0.772 0.483 0.828 0.515
±0.133 ±0.099 ±0.065 ±0.020 ±0.118 ±0.073 ±0.087 ±0.103 ±0.043 ±0.101
All Data 0.834 0.554 0.859 0.552 0.819 0.570 0.928 0.880 0.941 0.902
±0.037 ±0.114 ±0.065 ±0.217 ±0.069 ±0.096 ±0.049 ±0.046 ±0.020 ±0.018
III-D Modality Retention Module
$\mathbf{S}_{\text{retention}}=S_{\mathrm{r},u}+S_{\mathrm{i},u}.$ (11)
$\mathbf{T}_{\text{retention}}=T_{\mathrm{r},u}+T_{\mathrm{i},u}.$ (13)
III-E Style Clustering Module
To facilitate the alignment of disease-relevant features across modalities, MIRROR introduces a set of learnable cluster centers, $\mathbf{P}\in\mathbb{R}^{D}$, which are normalized to prevent degeneracy. The latent representations $z_{S}$ and $z_{T}$ are softly assigned to these centers, producing assignment distributions $\mathbf{S}_{\text{cluster}}$ and $\mathbf{T}_{\text{cluster}}$, respectively. The alignment between these assignments is encouraged through a bidirectional KL divergence penalty:
ensuring mutual consistency between the two modalities.
Formally, let $z_{S}$ and $z_{T}$ be $d$-dimensional latent vectors such that:
$L_{\text{style clustering}}=L_{\text{style}}+L_{\text{cluster}},$ (18)
TABLE II: Cancer survival analysis on TCGA-BRCA and TCGA-NSCLC, TCGA-RCC, and TCGA-COADREAD.
@@ III-E Style Clustering Module
[…]
TABLE III: Metadata of TCGA cohorts.
Cohort Subtype Total Samples
TCGA-BRCA Invasive Ductal Carcinoma (IDC) 955
Invasive Lobular Carcinoma (ILC)
TCGA-NSCLC Lung Adenocarcinoma (LUAD) 1,053
Lung Squamous Cell Carcinoma (LUSC)
TCGA-RCC Kidney Renal Clear Cell Carcinoma (KIRC) 943
Kidney Renal Papillary Cell Carcinoma (KIRP)
Kidney Chromophobe (KICH)
TCGA-COADREAD Colon Adenocarcinoma (COAD) 623
Rectum Adenocarcinoma (READ)
@@ III-F Global Optimization Objective
$L=\lambda_{\alpha}L_{\text{align}}+\lambda_{\beta}L_{\text{retention}}+\lambda_{\gamma}L_{\text{style clustering}},$ (19)
where $\lambda_{\alpha}$, $\lambda_{\beta}$, and $\lambda_{\gamma}$ are hyperparameters balancing each term. Minimizing this composite loss function promotes the enrichment of disease-relevant information, capturing both modality-shared and modality-specific components, denoted by $S_{r,s}$, $S_{r,u}$, $T_{r,s}$, and $T_{r,u}$. Simultaneously, it suppresses disease-irrelevant information, represented by $S_{i,s}$, $S_{i,u}$, $T_{i,s}$, and $T_{i,u}$. This selective optimization ensures that the learned feature representations concentrate on patterns essential for accurate disease characterization across modalities, thereby improving the model’s diagnostic and prognostic effectiveness.
Figure 4: Visualization of slide encoder attention weights on TCGA-BRCA, TCGA-NSCLC, TCGA-RCC and TCGA-COADREAD. Regions exhibiting higher attention scores predominantly correspond to malignant, tumor-bearing tissue, whereas areas with lower scores typically indicate normal regions.
@@ IV-A Datasets
Four distinct cohorts from the publicly available TCGA dataset [17] were utilized for evaluation: Breast Invasive Carcinoma (BRCA), Non-Small Cell Lung Cancer (NSCLC), Renal Cell Carcinoma (RCC), and Colon and Rectal Adenocarcinoma (COADREAD). Each cohort comprises specific subtypes, as detailed in Table III.
WSIs obtained from these cohorts were processed into patches at 20×20$\times$20 × magnification, and feature representations were extracted using ResNet-50 [53] and Phikon [54]. Corresponding transcriptomics data were collected from Xena [55] and preprocessed using the proposed novel pipeline, reducing the original 198,620 genes to 40,146 for the BRCA cohort, 10,234 for the NSCLC cohort, 10,303 for the RCC cohort, and 20,056 for the COADREAD cohort, forming the novel refined transcriptomics datasets.
@@ IV-B1 Data Preprocessing
For WSIs, adhering to the conventional preprocessing protocol established in [56], the foreground tissue regions are first segmented using the Otsu’s method [57]. These segmented regions are then divided into patches, forming an instance bag that is processed by a pre-trained patch encoder to extract feature representations. To enable batched training, we perform random sampling to select a fixed number of feature representations, employing replacement if the available number of representations is less than the required count.
@@ IV-C Downstream Tasks
Two tasks are used for downstream evaluations, namely subtype classification and survival analysis. We compared the proposed MIRROR against ABMIL [60], PORPOISE [61], Linear Classifier, and TANGLE [12], utilizing two distinct backbones under both 10-shot and linear probing settings. Among these methods, ABMIL and Linear Classifier take histopathology as input, whereas the other models integrate both histopathology and transcriptomics data. To ensure fair comparisons and optimal performance, all methods are trained on the proposed transcriptomic datasets. Notably, the original TANGLE framework evaluates only histopathology features. To comprehensively assess its performance and maintain methodological consistency, we extend TANGLE by incorporating both histopathology and transcriptomic features, aligning its input with that of MIRROR.
@@ IV-C1 Subtype Classification
As illustrated in Table I, the proposed MIRROR consistently achieves superior performance, demonstrating its effectiveness in integrating histopathology and transcriptomic data. On the TCGA-NSCLC dataset, MIRROR achieves the highest accuracy and F1 score across all settings, outperforming all baseline methods. Notably, its strong performance remains consistent regardless of the chosen backbone architecture, highlighting its robustness on the TCGA-NSCLC dataset. Similarly, on the TCGA-BRCA dataset, MIRROR achieves state-of-the-art results, surpassing the second best model’s (TANGLE) result by 1.5% in F1 score under linear probing setting with the Phikon backbone. Additionally, MIRROR significantly outperforms TANGLE by 6.4% in F1 score on the TCGA-RCC dataset under the 10-shot setting with the Phikon backbone. On the TCGA-COADREAD dataset, it surpasses TANGLE by 1.3% in accuracy under the linear probing setting with the Phikon backbone. These findings validate the robustness and generalizability of MIRROR across different cancer subtypes and backbone architectures.
@@ IV-D Ablation Study
TABLE IV: Ablation Study of Model Components on TCGA-BRCA.
@@
We conducted an ablation study to evaluate the effectiveness of the key components of MIRROR, including the modality alignment module, the modality retention module, and the style clustering module. The study was performed on the TCGA-BRCA dataset using linear probing with Phikon as the backbone, comparing MIRROR’s performance both with and without these components.
As shown in Table IV, the results highlight the contributions of each module. The modality retention module, while leading to a slight decline in survival analysis performance, enhances subtype classification, demonstrating its ability to preserve modality-specific information. Additionally, the introduction of the style clustering module encourages the model to focus on disease-relevant patterns while mitigating redundancy, enabling the extraction of more informative features. This refinement is reflected in the improved feature retention performance observed during training. With all modules incorporated, MIRROR achieves the best overall performance, confirming the synergistic benefits of the proposed components.
@@
As illustrated in Figure 4, we selected four representative samples from each cohort used, respectively. The attention weights from the slide encoder are visualized using heatmaps to highlight regions of interest. Areas shown in warmer colors indicate regions which the model pays more attention to. The results suggest MIRROR consistently attends to disease-relevant regions, particularly on cancerous regions. This strong alignment showcases MIRROR’s extraordinary interpretability, confirming that it not only excels in downstream tasks but also provides biologically grounded decision-making insights.
IV-E Qualitative Analysis
Figure 5: Visualization of histopathology and transcriptomics features encoded by MIRROR on the TCGA-NSCLC dataset, compared to those obtained using TANGLE. Pink dots represent samples from TCGA-LUSC, while blue dots represent samples from TCGA-LUAD. MIRROR clearly yields more distinct and representative feature distributions.
Additionally, as depicted in Figure 5, we visualize features extracted by encoders from TANGLE and MIRROR on the TCGA-NSCLC dataset. Pink and blue dots represent TCGA-LUSC and TCGA-LUAD samples, respectively. Features from MIRROR exhibit significantly improved class separability, with minimal overlap between subtypes. In contrast, TANGLE’s feature space demonstrates considerable inter-class mixing, suggesting weaker discriminative capability. The Euclidean distances between the two modalities are also computed for each method. Notably, the distance for MIRROR is reduced by 29.02% compared to TANGLE, indicating a substantially improved alignment between modalities. These findings highlight the advantage of MIRROR in generating more representative and well-aligned feature distributions, ultimately contributing to superior model performance on downstream tasks.